How many moles of PCl5 can be produced from 19.3 g of P4 (and excess Cl2)?
1
Expert's answer
2019-04-10T06:26:09-0400
"P_4 +10 Cl_2 \rightarrow 4PCl_5"

"n=\frac{m}{M}"

"n(P_4)=\frac{19.3 g}{123.9\frac{g}{mol}} = 0.156 mol"

As mole ratio "n(P_4):n(PCl_5)=1:4", then "n(PCl_5)= 4\times n(P_4)=4\times 0.156 mol = 0.623 mol"